I'm trying to find out how to do these problems examples are:
3n+5=6
4+5x=1
4y-15=-10
Please help

Answer by rfer(16322)

3n+5=6
subtract 5 from both sides
3n=1
divide both sides by 3
n=1/3
----------------------
4+5x=1
same method as above
-------------------------
4y-15=-10
add 15 to both sides
4y=-10+15
4y=5
divide by 4
y=5/4
or
y=1 1/4
whichever your teacher wants, both are correct